# Benchmarks for protocol control in nonequilibrium statistical mechanics

### v1

@@ I Introduction
This paper is the README file for a set of computer codes that simulate five models in the field of nonequilibrium statistical mechanics. Our aim in providing this code is to allow users to test methods designed to optimize time-dependent protocols. Recent work in this field has been driven by an interest in controlling strongly fluctuating experiments, and exploring the results of stochastic thermodynamics, including fluctuation relations, speed limits, and uncertainty relations Seifert (2012); Proesmans et al. (2020a); Blaber and Sivak (2023); Ciliberto (2017); Alvarado et al. (2025). A variety of numerical methods have been developed for these or closely-related purposes, including numerical methods Rotskoff and Crooks (2015) for evaluating thermodynamic geodesics Sivak and Crooks (2012); Zhong and DeWeese (2024), path-sampling methods Gingrich et al. (2016), variational methods Das and Limmer (2023), Monte Carlo methods Werner and Hartmann (2024), evolutionary methods Whitelam (2023a), and gradient-based methods Geiger and Dellago (2010); Rengifo et al. (2025); Engel et al. (2023).
@@
The problems described in this paper are intended for users to test and calibrate their protocol-control methods before deploying them on their problem of interest, motivated in part by the widespread use of benchmarks in machine learning. Our intent is to provide a set of problems that offer various tests of an optimization method. We have chosen problems that should be solvable, or nearly so, by a variety of methods.
@@
The problems were introduced by other authors, and have been selected for their physical interest and varied challenges. We consider overdamped Schmiedl and Seifert (2007) and underdamped Gomez-Marin et al. (2008) Langevin particles in a moving harmonic trap; information erasure using an underdamped Langevin particle Dago et al. (2021a, 2024); an active Brownian particle confined in a trap Baldovin et al. (2023); and magnetization reversal in the Ising model Rotskoff and Crooks (2015). In the trap problems the aim is to translate the trap with least work. In the erasure problem the aim is to perform erasure as reliably as possible. The active particle must be brought from a passive equilibrium state to a designated active steady state, while extracting from the active bath as much work as possible. The state of the Ising model must be changed with as little entropy produced as possible. All tasks must be accomplished in finite time $t_{\rm f}$, which is chosen to be short relative to the timescales of each system. This choice challenges analytic methods that identify efficient protocols in the long-time limit Sivak and Crooks (2012), and those analytic methods in turn provide valuable bounds with which to assess the effectiveness of numerical methods.
@@
The physical complexity of the systems increases steadily throughout the set, moving from a one-body system with one control parameter to a many-body system with two control parameters. However, this increase of system complexity does not necessarily correspond to the difficulty of controlling these systems as efficiently as possible. We have chosen problems whose difficulty is roughly comparable to that of MNIST mni in machine learning: these problems are simple enough to run on a laptop, accessible to a number of methods, but contain enough complexity to provide a non-trivial test of a control algorithm. One point to note: the success of a protocol is determined by its performance with respect to an order parameter of interest. For many problems it is possible to observe protocols that are visually quite different, but that perform similarly.
@@
In Section III we provide a worked example of protocol learning, providing C++ and Python implementations of a neurevolutionary optimization algorithm for the overdamped trap-translation problem, and a JAX implementation of automatic differentiation for the same. We also discuss the challenges associated with the other systems.
@@
The codes are available at the GitHub repository protocol-benchmarks/NESTbench25, in the folders c_code, python_code, and jax_code. Comments, suggestions, portability issues, and bug reports are welcome: please use the email addresses listed.
@@ II.2 Translating an underdamped particle with minimum work
in units of $k_{\rm B}T$. For our numerical implementation of the default (optimal) protocol, the function final_answer() returns an estimate of the mean work as order parameter $=2.90415\pm 0.0027$, using 100 samples of $10^{4}$ trajectories.
@@ II.3 Maximizing the probability of logic erasure using an underdamped particle
Here $S$ is the sign function, where $S(x)=-1$ (resp. 1) if $x<0$ (resp $x>0$). The fundamental lengthscale, timescale, and quality factor of the oscillator, $\sigma_{0}=1$ nm, $\omega_{0}=2\pi\times 1090$ Hz, and $Q=7$, are as in Section II.2. The integration scheme in the code engine_erasure is the same as in Section II.2.
@@
The function final_answer() computes the erasure failure rate over 100 samples of $10^{4}$ trajectories, and returns order parameter $=0.477974\pm 0.000456$ for the default trajectory. The goal of this problem is to find a protocol that makes this order parameter as small as possible.
@@
where $\bm{\xi}_{r}(\tau)$ and $\xi_{\theta}(\tau)$ are Gaussian white noise terms with zero mean and unit variance. The force term $-\kappa(t){\bm{r}}$ is derived from a harmonic restoring potential $U({\bm{r}},\kappa(t))=\frac{1}{2}\kappa(t){\bm{r}}^{2}$. The spring constant $\kappa(t)$ and swim speed $\lambda(t)$ are the control parameters ${\bm{c}}(t)$ of the problem.
@@ II.5 Minimizing entropy production during state change in the Ising model
The function load_default_protocol() calls the protocol of Ref. Rotskoff and Crooks (2015), which is the optimal protocol within the near-equilibrium approximation of Ref. Sivak and Crooks (2012). This protocol is shown in Fig. 5(a), with time as a parameter. This protocol avoids the first-order phase transition line and the critical point. With the default protocol, calling final_answer() gives an estimate for the entropy production of order parameter $=86.408\pm 0.0362$, using 100 samples of $10^{3}$ trajectories. To appreciate the scale of this result, note that simply switching to the final set of control parameters, without any change of configuration, produces entropy in the amount $2\times 32^{2}/0.7\approx 2.93\times 10^{3}$.
@@ III.1 Translating an overdamped particle with minimum work
The results of this learning process are shown in Fig. 6. Panel (a) shows the value of calculate_order_parameter(10000) for the best-performing member of the population of 50, as a function of generation number $n$. Panel (b) shows the protocol output_protocol() learned at the end of the process, which is similar to that of the optimal protocol shown in Fig. 3(a). It is not exactly the same, but its performance and that of the optimal protocol are difficult to distinguish when evaluated over a finite number of trajectories. For the learned protocol, the average final_answer() returns the estimate order parameter $=8.33697\pm 0.00439$, which is equal, within statistical error, to the optimal mean work value of $25/3$. Panel (c) shows the work distributions, generated by visualize_protocol(), for the learned and optimal protocols, which are essentially identical.
@@ III.5 Minimizing entropy production during state change in the Ising model
Fig. 10 shows the protocol learned by neuroevolution when instructed to change state in the Ising model while minimizing entropy production, overlaid on the default protocol (the near-equilibrium protocol of Ref. Rotskoff and Crooks (2015)). The learned protocol is markedly asymmetric, a feature shared by the protocol of Ref. Engel et al. (2023), which was identified by gradient descent. The function final_answer() shows the learned protocol to produce entropy in the amount order parameter $=52.6909\pm 0.030566$, about 61% of that produced by the near-equilibrium approximation.
@@ III.6 A note on genetic algorithms
trap_overdamped Overdamped Langevin particle in a moving harmonic trap Minimize work 8.34 $25/3$
trap_underdamped Underdamped Langevin particle in a moving harmonic trap Minimize work 3.14 2.91
erasure Logic erasure via underdamped dynamics Minimize error rate $1.4\times 10^{-3}$ Unknown
abp Active Brownian particle in a harmonic trap Enact state-to-state transformation and maximize extracted work $-3.0\times 10^{-3}$ ($\Delta=5.1\times 10^{-6}$, $\langle W\rangle=-2.9$) Unknown
ising 2D Ising model undergoing state change Minimize entropy production 52.7 Unknown
@@
IV Acknowledgements

### v2

# Benchmark control problems in nonequilibrium statistical mechanics

@@ I Introduction
Recent interest in finding optimal time-dependent protocols for stochastic dynamics has been motivated by the need to control strongly fluctuating experiments and the desire to explore the results of stochastic thermodynamics, including fluctuation relations, speed limits, and uncertainty relations  Seifert (2012); Proesmans et al. (2020a); Blaber and Sivak (2023); Ciliberto (2017); Alvarado et al. (2025). A variety of numerical approaches have been developed for these purposes, including methods for evaluating thermodynamic geodesics Rotskoff and Crooks (2015); Sivak and Crooks (2012); Zhong and DeWeese (2024), path-sampling techniques Gingrich et al. (2016), variational methods Das and Limmer (2023), Monte Carlo schemes Werner and Hartmann (2024), evolutionary strategies  Whitelam (2023a), and gradient-based algorithms Geiger and Dellago (2010); Rengifo et al. (2025); Engel et al. (2023).
@@
In this paper we introduce a set of benchmark problems and associated computer codes designed to enable users to test methods designed to optimize time-dependent protocols. This set of problems, motivated in part by the widespread use of benchmarks in machine learning, offers various tests of an optimization method, and should be solvable, or nearly so, by a variety of methods.
@@
The problems in this set were introduced by other authors, and we have selected them for their physical interest and varied challenges. We consider overdamped Schmiedl and Seifert (2007) and underdamped Gomez-Marin et al. (2008) Langevin particles in a moving harmonic trap; information erasure using an underdamped Langevin particle Dago et al. (2021a, 2024); an active Brownian particle confined in a trap Baldovin et al. (2023); and magnetization reversal in the Ising model Rotskoff and Crooks (2015). In the trap problems the aim is to translate the trap with least work. In the erasure problem the aim is to perform erasure as reliably as possible. The active particle must be brought from a passive equilibrium state to a designated active steady state, while extracting from the active bath as much work as possible. The state of the Ising model must be changed with as little entropy produced as possible. All tasks must be accomplished in finite time $t_{\rm f}$, which is chosen to be short relative to the timescales of each system. This choice challenges analytic methods that identify efficient protocols in the long-time limit Sivak and Crooks (2012), and those analytic methods in turn provide valuable bounds with which to assess the effectiveness of numerical methods.
@@
The physical complexity of the systems increases steadily throughout the set, moving from a one-body system with one control parameter to a many-body system with two control parameters. However, this increase of system complexity does not necessarily correspond to an increase in the difficulty of controlling these systems as efficiently as possible. We have chosen problems whose difficulty is roughly comparable to that of MNIST mni in machine learning: problems are simple enough to run on a laptop, accessible to a number of methods, but contain enough complexity to provide a non-trivial test of a control algorithm. One point to note: the success of a protocol is determined by its performance with respect to an order parameter of interest. For many problems it is possible to observe protocols that are visually quite different, but that perform similarly.
@@
In Section III we provide a worked example of protocol learning, providing C++ and Python implementations of a neurevolutionary learning algorithm for the overdamped trap-translation problem, and a JAX implementation of automatic differentiation for the same. We also provide the results of neurevolutionary learning for the other problems, and discuss some of the challenges associated with them.
@@
The codes are available at the GitHub repository protocol-benchmarks/NESTbench25, in the folders c_code, python_code, and jax_code. This paper serves as the README file for this set of codes. Comments, suggestions, portability issues, and bug reports are welcome: please use the email addresses listed.
@@ II.2 Translating an underdamped particle with minimum work
in units of $k_{\rm B}T$. For our numerical implementation of the default (optimal) protocol, the function final_answer() returns an estimate of the mean work as order parameter $2.90415\pm 0.0027$, using 100 samples of $10^{4}$ trajectories.
@@ II.3 Maximizing the probability of logic erasure using an underdamped particle
Here $S$ is the sign function, where S(x)=1S(x)=-1 (resp. 1) if $x<0$ (resp $x>0$). The fundamental lengthscale, timescale, and quality factor of the oscillator, $\sigma_{0}=1$ nm, $\omega_{0}=2\pi\times 1090$ Hz, and $Q=7$, are as in Section II.2. The integration scheme in the code engine_erasure is the same as in Section II.2.
@@
The function final_answer() computes the erasure failure rate over 100 samples of $10^{4}$ trajectories, and returns order parameter $0.477974\pm 0.000456$ for the default trajectory. The goal of this problem is to find a protocol that makes this order parameter as small as possible.
@@
where $\bm{\xi}_{r}(\tau)$ and $\xi_{\theta}(\tau)$ are Gaussian white noise terms with zero mean and unit variance. The force term $\kappa(t){\bm{r}}$ is derived from a harmonic restoring potential $U({\bm{r}},\kappa(t))=\frac{1}{2}\kappa(t){\bm{r}}^{2}$. The spring constant $\kappa(t)$ and swim speed $\lambda(t)$ are the control parameters ${\bm{c}}(t)$ of the problem.
@@ II.5 Minimizing entropy production during state change in the Ising model
The function load_default_protocol() calls the protocol of Ref. Rotskoff and Crooks (2015), which is the optimal protocol within the near-equilibrium approximation of Ref. Sivak and Crooks (2012). This protocol is shown in Fig. 5(a), with time as a parameter. This protocol avoids the first-order phase transition line and the critical point. With the default protocol, calling final_answer() gives an estimate for the entropy production of order parameter $86.408\pm 0.0362$, using 100 samples of $10^{3}$ trajectories. To appreciate the scale of this result, note that simply switching to the final set of control parameters, without any change of configuration, produces entropy in the amount $2\times 32^{2}/0.7\approx 2.93\times 10^{3}$.
@@ III.1 Translating an overdamped particle with minimum work
The results of this learning process are shown in Fig. 6. Panel (a) shows the value of calculate_order_parameter(10000) for the best-performing member of the population of 50, as a function of generation number $n$. Panel (b) shows the protocol output_protocol() learned at the end of the process, which is similar to that of the optimal protocol shown in Fig. 3(a). It is not exactly the same, but its performance and that of the optimal protocol are difficult to distinguish when evaluated over a finite number of trajectories. For the learned protocol, the average final_answer() returns the estimate order parameter $8.33697\pm 0.00439$, which is equal, within statistical error, to the optimal mean work value of $25/3$. Panel (c) shows the work distributions, generated by visualize_protocol(), for the learned and optimal protocols, which are essentially identical.
@@ III.5 Minimizing entropy production during state change in the Ising model
Fig. 10 shows the protocol learned by neuroevolution when instructed to change state in the Ising model while minimizing entropy production, overlaid on the default protocol (the near-equilibrium protocol of Ref. Rotskoff and Crooks (2015)). The learned protocol is markedly asymmetric, a feature shared by the protocol of Ref. Engel et al. (2023), which was identified by gradient descent. The function final_answer() shows the learned protocol to produce entropy in the amount order parameter $52.6909\pm 0.030566$, about 61% of that produced by the near-equilibrium approximation.
@@ III.6 A note on genetic algorithms
trap_overdamped Overdamped Langevin particle in a moving harmonic trap Minimize work $8.33697\pm 0.00439$ $25/3$
trap_underdamped Underdamped Langevin particle in a moving harmonic trap Minimize work $3.13804\pm 0.00283$ 2.91
erasure Logic erasure via underdamped dynamics Minimize error rate $1.446\times 10^{-3}\pm 3.57329\times 10^{-5}$ Unknown
abp Active Brownian particle in a harmonic trap Enact state-to-state transformation and maximize extracted work $3.0\times 10^{-3}$ ($\Delta=5.1\times 10^{-6}$, $\langle W\rangle=-2.9$) Unknown
ising 2D Ising model undergoing state change Minimize entropy production $52.6909\pm 0.030566$ Unknown
@@
IV Conclusions
We have presented a set of computer codes designed to test methods for optimizing time-dependent control protocols in fluctuating nonequilibrium systems. The problem set, available at the GitHub repository protocol-benchmarks/NESTbench25, contains five models, each with an associated optimization objective, coded in C++ and Python. These benchmark systems are simple enough to run on a laptop, but challenging enough to test the capabilities of modern optimization methods. We have included with the set a worked example and a set of results derived from neuroevolutionary learning, and we have discussed some of the challenges associated with each problem.
Benchmarks play an important role in the machine-learning literature by providing standardized test problems. Our benchmark set is designed to provide a set of standardized tests for physical scientists working on control problems. Our hope is that the set is instructive, and allows users to develop understanding of the strengths and weaknesses of particular methods of optimization.
V Acknowledgements